OpenStudy (nikki_skellington):

The main show tank has a radius of 60 feet and forms a quarter-sphere. What is the volume of the quarter-sphere shaped tank? Round your answer to the nearest whole number. You must explain your answer using words, and you must show all work and calculations to receive credit.

OpenStudy (3mar):

so what you feel?

OpenStudy (nikki_skellington):

im not sure....not good at this stuff

OpenStudy (3mar):

just choose and I will guide you In Sha' Allah

OpenStudy (nikki_skellington):

ok...but im unsure on what to do like where do i start?

OpenStudy (3mar):

We have a sphere ok? what is its volume? knowing just its radius!

OpenStudy (nikki_skellington):

60...?

OpenStudy (3mar):

what does 60 represent?

OpenStudy (nikki_skellington):

its volume :/

OpenStudy (3mar):

how is that? \[\frac{ 1 }{ 3 }\pi *r^3=\frac{ 1 }{ 3 }\pi *(60)^3=???\]

OpenStudy (3mar):

@Nikki_Skellington

OpenStudy (nikki_skellington):

sorry im trying to solve it...im not getting anything tho

OpenStudy (3mar):

One minute please

OpenStudy (nikki_skellington):

ok

OpenStudy (3mar):

are you stuck at the last equation I wrote?

OpenStudy (nikki_skellington):

yea

OpenStudy (3mar):

where?

OpenStudy (3mar):

What is not logic for you?

OpenStudy (nikki_skellington):

honestly, all of it im really not good at this type of math

OpenStudy (3mar):

Can you follow step by step?

OpenStudy (3mar):

The volume of a sphere is\[\frac{ 4 }{ 3 }\pi r^3\] we have -in this case - 1/4 sphere, not a complete sphere so its volume will be \[\frac{ 1 }{ 3 }\pi r^3\]

OpenStudy (3mar):

For the given data, the radius of this sphere is 60 ft (seems too big) just substitute with r in the last equation I wrote.

OpenStudy (nikki_skellington):

would it be 72,000?

OpenStudy (3mar):

you mean \[72000\pi\]

OpenStudy (nikki_skellington):

yup sorry

OpenStudy (3mar):

No problem.

OpenStudy (nikki_skellington):

so its correct?

OpenStudy (3mar):

Yes Congratulations!

OpenStudy (nikki_skellington):

yay! so thats it

OpenStudy (3mar):

Thank you for learning!!

OpenStudy (nikki_skellington):

no, thank you :)

OpenStudy (3mar):

I appreciate!

OpenStudy (3mar):

Any more questions?

OpenStudy (nikki_skellington):

i actually do have 1 more wait one sec please.

OpenStudy (3mar):

Ok

OpenStudy (nikki_skellington):

The holding tanks are congruent in size, and both are in the shape of a cylinder that has been cut in half vertically. The bottom of the tank is a curved surface. What is the volume of both tanks if the radius of tank #1 is 30 feet and the height of tank #2 is 110 feet? You must explain your answer using words, and you must show all work and calculations to receive credit.

OpenStudy (3mar):

Can we start in new post, please?

OpenStudy (nikki_skellington):

yea

OpenStudy (3mar):

thanks